OpenStudy (anonymous):

Use the three steps to solve the problem. One of two complementary angles is 4 more than the other. Find the angles. (Recall that complementary angles are angles whose sum is equal to 90 .)

ganeshie8 (ganeshie8):

complementary angles add up to 90

ganeshie8 (ganeshie8):

so, if we let one angle =x, other angle = x+4

ganeshie8 (ganeshie8):

step 1 : solve for x

OpenStudy (anonymous):

86

ganeshie8 (ganeshie8):

correct ! x = 86

ganeshie8 (ganeshie8):

wait a sec, did we get x=86 ?

ganeshie8 (ganeshie8):

x + x + 4 = 90 2x + 4 = 90 2x = 86 x = ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

x + x + 4 = 90 2x + 4 = 90 2x = 86 x = 86/2 = ?

OpenStudy (anonymous):

43

ganeshie8 (ganeshie8):

thanks god !

ganeshie8 (ganeshie8):

:)

ganeshie8 (ganeshie8):

thats step 1,

ganeshie8 (ganeshie8):

step 2 : first angle = x = 43

ganeshie8 (ganeshie8):

step 3 : second angle = x+4 = ?

OpenStudy (anonymous):

47

ganeshie8 (ganeshie8):

perfect ! good work !

OpenStudy (anonymous):

thanks

ganeshie8 (ganeshie8):

np :)